OpenStudy (anonymous):

Evaluate the integral \[\int\limits_{}^{}\int\limits_{}^{}R(x ^{2}-2y ^{2})dA\] where R is the first quadrant region between the circles of radius 3 and radius 6

OpenStudy (mathmate):

Starting with what you started before, ∫∫R(x2−2y2)dA =∫∫(rcos(t))^2-2(rsin(t))^2rdtdr =∫∫r^3(cos(t)^2-2sin(t)^2)dtdr =∫∫r^3(cos(t)^2-2sin(t)^2)dt =[r^4/4]∫((1/2)(1+cos(2t))-(1-cos(2t))dt =(6^4-3^4)/4 * ∫(-1/2 + cos(2t)/2 + cos(2t))dt =(1215/4) [-t/2 + sin(2t)/4 + sin(2t)/2] from 0 to pi/2 =(1215/4)(-pi/4) =-1215pi/16 I believe the negative value is from the -2 in -2y^2.